# Gauge freedoms in unravelled quantum dynamics: When do different continuous measurements yield identical quantum trajectories?

### v1

@@ V.4 Permutations
V.5 General representations
Finally, we observe that the representation in Eq. (V.3) is actually a minimal representation of the QME. Therefore, a general representation of this QME can be obtained by the transformation Eq. (4), which shifts and mixes the jumps, together with an appropriate change in the Hamiltonian. The transformation in Eq. (50) features no shifts, and it also has a block structure, as encoded by Eq. (31): this ensures that $H,J_{1},\dots,J_{5}$ and $H,J_{1}^{\prime},\dots,J_{3}^{\prime}$ lead to identical quantum trajectories, and equivalent partially-labelled quantum dynamics. In order to ensure equivalent (fully)-labelled quantum dynamics, the isometric mixing of jump operators must reduce to simple permutation and multiplication by phases as in Eq. (48).
VI Proof of Theorem 1
In this section, we give the proof of Theorem 1, which was presented in Sec. III.
VI.1 Conditions for equality of unravelled generators
@@
In summary, we have shown the two conditions in Eqs. (58) and (60) are together equivalent to Eq. (51), which in turn means that the unravelled generators are equal for the two representations. We refer to Eqs. (58) and (60) as the drift condition and the jump condition, respectively. We analyse these conditions separately, before combining the results in Sec. VI.4 to prove Theorem 1.
VI.2 Jump condition
@@
VI.2.1 A set of states where every SJED has a distinct destination
@@
To this end, write ${\cal D}_{k}(\mathbf{C})=\{\mathcal{D}_{k}(\psi):\psi\in\mathbf{C}\}$ for the set of destinations from $\mathbf{C}$ facilitated by jump operator $J_{k}$, and similarly $\tilde{\cal D}_{j}(\mathbf{C})$ for the analogous set facilitated by jump operator $\tilde{J}_{j}$, see Fig. 2. To show that (i) holds throughout $\mathbf{C}$, note that the rates $r_{k}(\psi_{0})$, $\tilde{r}_{j}(\psi_{0})$ are non-zero for all $j$ and $k$: hence by linearity [cf. Eq. (10)] there exist sufficiently small $\delta>0$ such that (i) follows for all $\psi\in\mathbf{C}$ [cf. Eq. (11)]. Similarly, (ii) holds for $\psi=\psi_{0}$ which means that the destinations ${\cal D}_{k}(\psi_{0})$ are different for each SJED. Furthermore, the destinations are continuous functions of $\psi$ provided that their rate are non-zero, which is already guaranteed by (i). Thus there exists $\delta>0$ [as chosen for (i) or smaller] such that ${\cal D}_{k}(\mathbf{C})$ is disjoint from ${\cal D}_{k^{\prime}}(\mathbf{C})$ if $k\in S_{\alpha}$ and $k^{\prime}\in S_{\alpha^{\prime}}$ with $\alpha\neq\alpha^{\prime}$, see Fig. LABEL:fig:sketch_a. Then indeed the property (ii) is true for all $\psi\in\mathbf{C}$. The argument for property (iii) is exactly analogous.
@@
VI.2.2 Singleton SJEDs
From the jump condition expressed as Eq. (61), we can expect that the action of SJEDs must lead to the same sets of destinations for both representations. In this section, we assume that all SJEDs are singletons (i.e., each SJED contains a single jump). Then, in the setting of Fig. LABEL:fig:sketch_a, this requires that the destination sets $\mathcal{D}_{1}(\mathbf{C}),\dots,{\mathcal{D}}_{d}(\mathbf{C})$ should be in one-to-one correspondence with $\tilde{\mathcal{D}}_{1}(\mathbf{C}),\dots,\tilde{\mathcal{D}}_{\tilde{d}}(\mathbf{C})$. Formalising this observation, we derive conditions on the jump operators such that Eq. (61) is satisfied [see Eq. (69) below]. This is helpful to build physical intuition before discussing the case of general SJEDs in Sec. VI.2.3.
@@ VI.2.2 Singleton SJEDs
with $c\in\mathbb{C}$ and $z_{c}$ a normalisation constant. Recall that the set $\mathbf{C}$ is an intersection of pure states within a neighbourhood of $\psi_{0}$, so for any $|\Delta\rangle$ there exists a finite range of $c$, including $c=0$, such that $\psi_{c}\in\mathbf{C}$. We consider Eq. (68) for $\psi=\psi_{0}$ and $\psi=\psi_{c}$, and multiply by $z_{0}$ and $z_{c}$, respectively, to get
@@
$\langle\psi_{0}|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\psi_{0}\rangle$ (73)
$\quad+c\,\langle\psi_{0}|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\Delta\rangle$
$\quad+c^{*}\langle\Delta|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\psi_{0}\rangle$
$\quad+|c|^{2}\langle\Delta|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\Delta\rangle=0$
@@
$\langle\psi_{0}|\tilde{J}_{k}^{\dagger}\tilde{J}_{k}|\psi_{0}\rangle+c\langle\psi_{0}|\tilde{J}_{k}^{\dagger}\tilde{J}_{k}|\Delta\rangle$ (76)
$={\rm e}^{i(\phi_{c}-\phi_{0})}\left[\langle\psi_{0}|{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}|\psi_{0}\rangle+c\langle\psi_{0}|{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}|\Delta\rangle\right],$
@@
VI.2.3 General SJEDs
We now return to general conditions that are necessary and sufficient for Eq. (60) to be valid. These are needed for the proof of Theorem 1. The reasoning is analogous to Sec. VI.2.2.
@@ VI.2.3 General SJEDs
$\sum_{\beta=1}^{\tilde{d}_{C}}\delta\!\left\{\psi^{\prime}-\frac{\tilde{\mathcal{A}}_{\beta}(\psi)}{\text{Tr}[\tilde{\mathcal{A}}_{\beta}(\psi)]}\right\}\text{Tr}[\tilde{\mathcal{A}}_{\beta}(\psi)]\\ =\sum_{\alpha=1}^{d_{C}}\delta\!\left\{\psi^{\prime}-\frac{\mathcal{A}_{\alpha}(\psi)}{\text{Tr}[\mathcal{A}_{\alpha}(\psi)]}\right\}\text{Tr}[\mathcal{A}_{\alpha}(\psi)].$ (79)
@@
$\tilde{\cal A}_{\alpha}(\psi)={\cal A}_{\pi(\alpha)}(\psi),\quad\forall\alpha=1,...,d_{C}\quad\forall\psi\in\mathbf{C}\,,$ (81)
where $\pi$ is a permutation of $\{1,2,\dots,d_{C}\}$ independent of $\psi\in\mathbf{C}$.
@@
$\tilde{\cal A}_{\alpha}={\cal A}_{\pi(\alpha)},\quad\forall\alpha=1,...,d_{C},$ (82)
@@
From Eq. (81) we see that if SJED $\tilde{S}_{\alpha}$ is of reset type, then $S_{\pi(\alpha)}$ must also be of reset type (and likewise for SJEDs of non-reset type), as the dimension of the image of $\mathcal{A}_{\alpha}(\psi)$ is $1$ for reset SJEDs and $>1$ for non-reset SJEDs. Hence, if $\alpha$ labels a reset-type SJED, then for $j\in\tilde{S}_{\alpha}$, we have $\tilde{J}_{j}=\sqrt{\tilde{\gamma}_{j}}|\tilde{\chi}_{\alpha}\rangle\!\langle\tilde{\xi}_{j}|$ [cf. Eq. (21)] and from Eq. (81)
$|\tilde{\chi}_{\alpha}\rangle\!\langle\tilde{\chi}_{\alpha}|=|\chi_{\pi(\alpha)}\rangle\!\langle\chi_{\pi(\alpha)}|.$ (83)
@@
$\text{Tr}(\tilde{\Gamma}_{\alpha}\psi)=\text{Tr}[\Gamma_{\pi(\alpha)}\psi].$ (84)
@@
$c\,\langle\psi_{0}|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi(\alpha)}\big{]}|\Delta\rangle+c^{*}\langle\Delta|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi(\alpha)}\big{]}|\psi_{0}\rangle\\ +|c|^{2}\langle\Delta|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi(\alpha)}\big{]}|\Delta\rangle=0.$ (85)
Analogously to Eq. (73), since this holds for complex $c$ in a finite neighbourhood of $c=0$ the coefficients for $c,c^{*},|c|^{2}$ need to match. In particular, from the term with $|c|^{2}$, by observing that $|\Delta\rangle$ is arbitrary and using Hermiticity of $\tilde{\Gamma}_{\alpha}-\Gamma_{\pi(\alpha)}$, we obtain
$\tilde{\Gamma}_{\alpha}=\Gamma_{\pi(\alpha)},$ (86)
@@
$|\tilde{\lambda}^{(\alpha)}|^{2}\tilde{\mathcal{J}}^{(\alpha)}(\psi)=|\lambda^{[\pi(\alpha)]}|^{2}\mathcal{J}^{[\pi(\alpha)]}(\psi),$ (87)
@@
$\tilde{\lambda}^{(\alpha)}\tilde{J}^{(\alpha)}|\psi\rangle={\rm e}^{i\phi_{\alpha\psi}}\lambda^{[\pi(\alpha)]}J^{[\pi(\alpha)]}|\psi\rangle,$ (88)
@@
$\tilde{\lambda}^{(\alpha)}\tilde{J}^{(\alpha)}={\rm e}^{i\phi_{\alpha}}\lambda^{[\pi(\alpha)]}J^{[\pi(\alpha)]}$ (89)
@@ VI.3 Drift term
To show first Eq. (90), we note that by definition of ${\cal B}(\psi)$ and ~$\tilde{\cal B}(\psi)$ [cf. Eq. (6)]
@@
with $V=\tilde{H}_{\text{eff}}-H_{\text{eff}}$ for all $\psi$. This implies Eq. (90) as follows. Let $|a\rangle\neq|b\rangle$ be two elements of an orthonormal basis for the system. Take $\psi=|a\rangle\!\langle a|$, and multiply Eq. (90) from the left by $\langle b|$ and from the right by $|a\rangle$. As $\langle a|b\rangle=0$, Eq. (90) then yields (from its second term) $\langle b|V|a\rangle=0$. As this holds for all pair of elements of the basis, $V$ is diagonal. Furthermore, since this holds for any orthogonal basis, we have
@@ VI.4 Final result
VII Proofs of Theorems 2 and 3
This section outlines the proofs of Theorems 2 and 3. They have analogous structure to the proof of Theorem 1 in Sec. VI.
VII.1 Proof of Theorem 2
@@
Taking $f(\psi,\bm{q})=g(\psi)$, we obtain Eq. (51), i.e., the gauge invariance for the unlabelled dynamics as $\Pi$ and $\Pi^{\dagger}$ acts on $\bm{q}$ only, which condition is equivalent to the drift condition in Eq. (58) and the jump condition in Eq. (60). From Sec. VI, we have already have that the condition in Eq. (28) in Theorem 1 is necessary for this. However, the conditions for Theorem 2 are stronger, and we prove below that they are indeed necessary.
@@
$0=\left[\tilde{\mathcal{B}}(\psi)-{\cal B}(\psi)\right]\cdot\nabla f(\psi,\bm{q})\\ +\sum_{k}\int d\psi^{\prime}\left\{\tilde{w}_{k}(\psi,\psi^{\prime})\left[f(\psi^{\prime},\bm{q}+\bm{e}_{\pi(k)})-f(\psi,\bm{q})\right]\right.\\ -w_{k}(\psi,\psi^{\prime})\left[f(\psi^{\prime},\bm{q}+\bm{e}_{k})-f(\psi,\bm{q})\right]\big{\}}.$ (98)
@@
for $k=1,\dots,d$ and all $\psi,\psi^{\prime}$. Note that the permutation $\pi$ in Eq. (102) is determined by $\Pi^{\dagger}$ in Eq. (97) [cf. Eq. (34)] (and not state dependent). Recalling the arguments of Sec. VI.2.2 of the Theorem 1 proof for the case of singleton SJEDs, one sees that the labelled jump condition in Eq. (103) requires both Eq. (64) and (65) to hold [together with Eq. (66)], for all $\psi$. Hence, Eq. (67) holds for any $\psi$, and we obtain [recall Eq. (29)]

### v3

@@ 5.4 Permutations
5.5 General representations
Finally, we observe that the representation in Eq. (49) is actually a minimal representation of the QME. Therefore, a general representation of this QME can be obtained by the transformation Eq. (4), which shifts and mixes the jumps, together with an appropriate change in the Hamiltonian. The transformation in Eq. (50) features no shifts, and it also has a block structure, as encoded by Eq. (31): this ensures that $H,J_{1},\dots,J_{5}$ and $H,J_{1}^{\prime},\dots,J_{3}^{\prime}$ lead to identical quantum trajectories, and equivalent partially-labelled quantum dynamics. In order to ensure equivalent (fully)-labelled quantum dynamics, the isometric mixing of jump operators must reduce to simple permutation and multiplication by phases as in Eq. (48).
6 Proof of Theorem 1
In this section, we give the proof of Theorem 1, which was presented in Sec. 3.
6.1 Conditions for equality of unravelled generators
@@
In summary, we have shown the two conditions in Eqs. (58) and (60) are together equivalent to Eq. (51), which in turn means that the unravelled generators are equal for the two representations. We refer to Eqs. (58) and (60) as the drift condition and the jump condition, respectively. We analyse these conditions separately, before combining the results in Sec. 6.4 to prove Theorem 1.
6.2 Jump condition
@@
6.2.1 A set of states where every SJED has a distinct destination
@@
To this end, write ${\cal D}_{k}(\mathbf{C})=\{\mathcal{D}_{k}(\psi):\psi\in\mathbf{C}\}$ for the set of destinations from $\mathbf{C}$ facilitated by jump operator $J_{k}$, and similarly $\tilde{\cal D}_{j}(\mathbf{C})$ for the analogous set facilitated by jump operator $\tilde{J}_{j}$, see Fig. 2. To show that (i) holds throughout $\mathbf{C}$, note that the rates $r_{k}(\psi_{0})$, $\tilde{r}_{j}(\psi_{0})$ are non-zero for all $j$ and $k$: hence by linearity [cf. Eq. (10)] there exist sufficiently small $\delta>0$ such that (i) follows for all $\psi\in\mathbf{C}$ [cf. Eq. (11)]. Similarly, (ii) holds for $\psi=\psi_{0}$ which means that the destinations ${\cal D}_{k}(\psi_{0})$ are different for each SJED. Furthermore, the destinations are continuous functions of $\psi$ provided that their rates are non-zero, which is already guaranteed by (i). Thus there exists $\delta>0$ [as chosen for (i) or smaller] such that ${\cal D}_{k}(\mathbf{C})$ is disjoint from ${\cal D}_{k^{\prime}}(\mathbf{C})$ if $k\in S_{\alpha}$ and $k^{\prime}\in S_{\alpha^{\prime}}$ with $\alpha\neq\alpha^{\prime}$, see Fig. LABEL:fig:sketch_a. Then indeed the property (ii) is true for all $\psi\in\mathbf{C}$. The argument for property (iii) is exactly analogous.
@@
6.2.2 Singleton SJEDs
From the jump condition expressed as Eq. (61), we can expect that the action of SJEDs must lead to the same sets of destinations for both representations. In this section, we assume that all SJEDs are singletons (i.e., each SJED contains a single jump). Then, in the setting of Fig. LABEL:fig:sketch_a, this requires that the destination sets $\mathcal{D}_{1}(\mathbf{C}),\dots,{\mathcal{D}}_{d}(\mathbf{C})$ should be in one-to-one correspondence with $\tilde{\mathcal{D}}_{1}(\mathbf{C}),\dots,\tilde{\mathcal{D}}_{\tilde{d}}(\mathbf{C})$. Formalising this observation, we derive conditions on the jump operators such that Eq. (61) is satisfied [see Eq. (69) below]. This is helpful to build physical intuition before discussing the case of general SJEDs in Sec. 6.2.3.
@@ 6.2.2 Singleton SJEDs
with $c\in\mathbb{C}$ and $z_{c}$ a normalisation constant. Recall that the set $\mathbf{C}$ is an intersection of pure states with a neighbourhood of $\psi_{0}$, so for any $|\Delta\rangle$ there exists a finite range of $c$, including $c=0$, such that $\psi_{c}\in\mathbf{C}$. We consider Eq. (68) for $\psi=\psi_{0}$ and $\psi=\psi_{c}$, and multiply by $z_{0}$ and $z_{c}$, respectively, to get
@@
$\langle\psi_{0}|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\psi_{0}\rangle+c\,\langle\psi_{0}|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\Delta\rangle\\ +c^{*}\langle\Delta|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\psi_{0}\rangle+|c|^{2}\langle\Delta|\big{[}\tilde{J}_{k}^{\dagger}\tilde{J}_{k}-{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}]|\Delta\rangle=0$ (73)
@@
$\langle\psi_{0}|\tilde{J}_{k}^{\dagger}\tilde{J}_{k}|\psi_{0}\rangle+c\langle\psi_{0}|\tilde{J}_{k}^{\dagger}\tilde{J}_{k}|\Delta\rangle={\rm e}^{i(\phi_{c}-\phi_{0})}\left[\langle\psi_{0}|{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}|\psi_{0}\rangle+c\langle\psi_{0}|{J}_{\pi(k)}^{\dagger}{J}_{\pi(k)}|\Delta\rangle\right],$ (76)
@@
6.2.3 General SJEDs
We now return to general conditions that are necessary and sufficient for Eq. (60) to be valid. These are needed for the proof of Theorem 1. The reasoning is analogous to Sec. 6.2.2.
@@ 6.2.3 General SJEDs
$\sum_{\beta=1}^{\tilde{d}_{C}}\delta\!\left\{\psi^{\prime}-\frac{\tilde{\mathcal{A}}_{\beta}(\psi)}{\text{Tr}[\tilde{\mathcal{A}}_{\beta}(\psi)]}\right\}\text{Tr}[\tilde{\mathcal{A}}_{\beta}(\psi)]=\sum_{\alpha=1}^{d_{C}}\delta\!\left\{\psi^{\prime}-\frac{\mathcal{A}_{\alpha}(\psi)}{\text{Tr}[\mathcal{A}_{\alpha}(\psi)]}\right\}\text{Tr}[\mathcal{A}_{\alpha}(\psi)].$ (79)
@@
$\tilde{\cal A}_{\alpha}(\psi)={\cal A}_{\pi_{\textsc{c}}(\alpha)}(\psi),\quad\forall\alpha=1,...,d_{C}\quad\forall\psi\in\mathbf{C}\,,$ (81)
where $\pi_{\textsc{c}}$ is a permutation of $\{1,2,\dots,d_{C}\}$ independent of $\psi\in\mathbf{C}$.
@@
$\tilde{\cal A}_{\alpha}={\cal A}_{\pi_{\textsc{c}}(\alpha)},\quad\forall\alpha=1,...,d_{C},$ (82)
@@
From Eq. (81) we see that if SJED $\tilde{S}_{\alpha}$ is of reset type, then $S_{\pi_{\textsc{c}}(\alpha)}$ must also be of reset type (and likewise for SJEDs of non-reset type), as the dimension of the image of $\mathcal{A}_{\alpha}(\psi)$ is $1$ for reset SJEDs and $>1$ for non-reset SJEDs. Hence, if $\alpha$ labels a reset-type SJED, then for $j\in\tilde{S}_{\alpha}$, we have $\tilde{J}_{j}=\sqrt{\tilde{\gamma}_{j}}|\tilde{\chi}_{\alpha}\rangle\!\langle\tilde{\xi}_{j}|$ [cf. Eq. (21)] and from Eq. (81)
$|\tilde{\chi}_{\alpha}\rangle\!\langle\tilde{\chi}_{\alpha}|=|\chi_{\pi_{\textsc{c}}(\alpha)}\rangle\!\langle\chi_{\pi_{\textsc{c}}(\alpha)}|.$ (83)
@@
$\text{Tr}(\tilde{\Gamma}_{\alpha}\psi)=\text{Tr}[\Gamma_{\pi_{\textsc{c}}(\alpha)}\psi].$ (84)
@@
$c\,\langle\psi_{0}|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi_{\textsc{c}}(\alpha)}\big{]}|\Delta\rangle+c^{*}\langle\Delta|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi_{\textsc{c}}(\alpha)}\big{]}|\psi_{0}\rangle+|c|^{2}\langle\Delta|\big{[}\tilde{\Gamma}_{\alpha}-\Gamma_{\pi_{\textsc{c}}(\alpha)}\big{]}|\Delta\rangle=0.$ (85)
Analogously to Eq. (73), since this holds for complex $c$ in a finite neighbourhood of $c=0$ the coefficients for $c,c^{*},|c|^{2}$ need to match. In particular, from the term with $|c|^{2}$, by observing that $|\Delta\rangle$ is arbitrary and using Hermiticity of $\tilde{\Gamma}_{\alpha}-\Gamma_{\pi_{\textsc{c}}(\alpha)}$, we obtain
$\tilde{\Gamma}_{\alpha}=\Gamma_{\pi_{\textsc{c}}(\alpha)},$ (86)
@@
$|\tilde{\lambda}^{(\alpha)}|^{2}\tilde{\mathcal{J}}^{(\alpha)}(\psi)=|\lambda^{[\pi_{\textsc{c}}(\alpha)]}|^{2}\mathcal{J}^{[\pi_{\textsc{c}}(\alpha)]}(\psi),$ (87)
@@
$\tilde{\lambda}^{(\alpha)}\tilde{J}^{(\alpha)}|\psi\rangle={\rm e}^{i\phi_{\alpha\psi}}\lambda^{[\pi_{\textsc{c}}(\alpha)]}J^{[\pi_{\textsc{c}}(\alpha)]}|\psi\rangle,$ (88)
@@
$\tilde{\lambda}^{(\alpha)}\tilde{J}^{(\alpha)}={\rm e}^{i\phi_{\alpha}}\lambda^{[\pi_{\textsc{c}}(\alpha)]}J^{[\pi_{\textsc{c}}(\alpha)]}$ (89)
@@ 6.3 Drift term
To show first Eq. (90), we use Eq. (58) with the definitions of ${\cal B}(\psi)$ and ~$\tilde{\cal B}(\psi)$ [cf. Eq. (6)] to obtain that
@@
for all $\psi$, where $V=\tilde{H}_{\text{eff}}-H_{\text{eff}}$. Let $|a\rangle\neq|b\rangle$ be two elements of an orthonormal basis for the system. Take $\psi=|a\rangle\!\langle a|$, and multiply Eq. (92) from the left by $\langle b|$ and from the right by $|a\rangle$. As $\langle a|b\rangle=0$, Eq. (92) then yields (from its first term) $\langle b|V|a\rangle=0$. As this holds for any pair of elements of the basis, $V$ is diagonal. Furthermore, since this holds for any orthogonal basis, we have
@@ 6.4 Final result
7 Proofs of Theorems 2 and 3
This section outlines the proofs of Theorems 2 and 3. They have analogous structure to the proof of Theorem 1 in Sec. 6.
7.1 Proof of Theorem 2
@@
Taking $f(\psi,\bm{q})=g(\psi)$, we obtain Eq. (51), i.e., the gauge invariance for the unlabelled dynamics as $\Pi$ and $\Pi^{\dagger}$ acts on $\bm{q}$ only, which condition is equivalent to the drift condition in Eq. (58) and the jump condition in Eq. (60). From Sec. 6, we have already have that the condition in Eq. (28) in Theorem 1 is necessary for this. However, the conditions for Theorem 2 are stronger, and we prove below that they are indeed necessary.
@@
$0=\left[\tilde{\mathcal{B}}(\psi)-{\cal B}(\psi)\right]\cdot\nabla f(\psi,\bm{q})+\sum_{k}\int d\psi^{\prime}\Big{\{}\tilde{w}_{k}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{q}+\bm{e}_{\pi(k)})-f(\psi,\bm{q})\big{]}\\ -w_{k}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{q}+\bm{e}_{k})-f(\psi,\bm{q})\big{]}\Big{\}}.$ (98)
@@
for $k=1,\dots,d$ and all $\psi,\psi^{\prime}$. Note that the permutation $\pi$ in Eq. (102) is determined by $\Pi^{\dagger}$ in Eq. (97) [cf. Eq. (34)] (and not state dependent). Recalling the arguments of Sec. 6.2.2 one sees that the labelled jump condition in Eq. (103) requires both Eq. (64) and (65) to hold [together with Eq. (66)], for all $\psi$. Hence, Eq. (67) holds for any $\psi$, and we obtain [recall Eq. (29)]